# Enhancing Fairness in Unsupervised Graph Anomaly Detection through Disentanglement

### v1

@@ A.1. Baselines
    (20) $\mathcal{L}_{DP}^{FairOD}=\left|\frac{\left(\sum_{i=1}^{n}{o}_{i}-\mu_{o}\right)\left(\sum_{i=1}^{n}{s}_{i}-\mu_{s}\right)}{\sigma_{o}\sigma_{S}}\right|$
@@
    (21) $\mathcal{L}_{ADCG}^{FairOD}=\sum_{s\in\{0,1\}}\left(1-\sum_{\{v_{i}:s_{i}=s\}}\frac{2^{{o}_{i}^{base}}-1}{\mathrm{DNM}_{s}}\right),$
@@
    (22) $\mathcal{L}^{Corr}=\left|\frac{\mathbf{o}\cdot\mathbf{S}}{\sqrt{(\mathbf{o}\cdot\mathbf{o})(\mathbf{S}\cdot\mathbf{S})}}\right|,$
@@
    (23) $\mathcal{L}_{DP}^{HIN}=\sum_{k\in\{0,1\}}\Bigg{(}\frac{\sum_{\{v_{i}:s_{i}=1\}}P(\hat{y}_{i}=k)}{|{\{v_{i}:s_{i}=1\}}|}-\frac{\sum_{\{v_{i}:s_{i}=0\}}P(\hat{y}_{i}=k)}{|{\{v_{i}:s_{i}=0\}}|}\Bigg{)}^{2},$
@@
    (24) $\begin{split}\mathcal{L}_{EO}^{HIN}=\sum_{k\in\{0,1\}}\Bigg{(}&\frac{\sum_{\{v_{i}:s_{i}=1,y=k\}}Pr(\hat{y}_{i}=k)}{|{\{v_{i}:s_{i}=1,y=k\}}|}\\ &-\frac{\sum_{\{v_{i}:s_{i}=0,y=k\}}Pr(\hat{y}_{i}=k)}{|{\{v_{i}:s_{i}=0,y=k\}}|}\Bigg{)}^{2}.\end{split}$
@@
    (25) $\mathcal{L}=\mathcal{L}_{base}+\lambda\mathcal{L}_{DP}^{HIN}+\gamma\mathcal{L}_{ADCG}^{FairOD}.$

### v2

@@ Appendix D Baselines
  • VGOD (Huang et al., 2023) proposes a variance-based framework that combines a variance-based model for structural outlier detection with an attribute reconstruction model for contextual outlier detection. It achieves balanced detection performance between structural and contextual outliers while addressing data leakage issues present in existing injection-based approaches.
@@
    $\mathcal{L}_{DP}^{FairOD}=\left|\frac{\left(\sum_{i=1}^{n}{o}_{i}-\mu_{o}\right)\left(\sum_{i=1}^{n}{s}_{i}-\mu_{s}\right)}{\sigma_{o}\sigma_{S}}\right|$ (20)
@@
    $\mathcal{L}_{ADCG}^{FairOD}=\sum_{s\in\{0,1\}}\left(1-\sum_{\{v_{i}:s_{i}=s\}}\frac{2^{{o}_{i}^{base}}-1}{\mathrm{DNM}_{s}}\right),$ (21)
@@
    $\mathcal{L}^{Corr}=\left|\frac{\mathbf{o}\cdot\mathbf{S}}{\sqrt{(\mathbf{o}\cdot\mathbf{o})(\mathbf{S}\cdot\mathbf{S})}}\right|,$ (22)
@@
    $\mathcal{L}_{DP}^{HIN}=\sum_{k\in\{0,1\}}\Bigg{(}\frac{\sum_{\{v_{i}:s_{i}=1\}}P(\hat{y}_{i}=k)}{|{\{v_{i}:s_{i}=1\}}|}-\frac{\sum_{\{v_{i}:s_{i}=0\}}P(\hat{y}_{i}=k)}{|{\{v_{i}:s_{i}=0\}}|}\Bigg{)}^{2},$ (23)
@@
    $\mathcal{L}_{EO}^{HIN}=\sum_{k\in\{0,1\}}\Bigg{(}\frac{\sum_{\{v_{i}:s_{i}=1,y=k\}}Pr(\hat{y}_{i}=k)}{|{\{v_{i}:s_{i}=1,y=k\}}|}-\frac{\sum_{\{v_{i}:s_{i}=0,y=k\}}Pr(\hat{y}_{i}=k)}{|{\{v_{i}:s_{i}=0,y=k\}}|}\Bigg{)}^{2}.$ (24)
@@
    $\mathcal{L}=\mathcal{L}_{base}+\lambda\mathcal{L}_{DP}^{HIN}+\gamma\mathcal{L}_{ADCG}^{FairOD}.$ (25)
@@
Appendix E Implementation Details
For DOMINANT, CoLA, CONAD and VGOD, we use the code and default hyper-parameters provided by PyGOD [^1] https://pygod.org (Liu et al., 2022c). For FairWalk [^2] https://github.com/urielsinger/fairwalk and EDITS [^3] https://github.com/yushundong/EDITS, we implement them using the code published by their authors. For FairOD, HIN, and Correlation, we implement the code provided by FairGAD [^4] https://github.com/nigelnnk/FairGAD.
Appendix F Limitations and Future Work
First, DEFEND encounters difficulties when protected attributes are highly correlated with other features, potentially resulting in information loss during disentanglement. Additionally, DEFEND relies on adversarial learning to approximate the Total Correlation penalty used for disentanglement. This dependency introduces potential convergence instability issues that might compromise the robustness of learned representations (Oh et al., 2022). We plan to extend the framework to handle multiple and continuous sensitive attributes simultaneously, which would enhance its practical applicability.